OpenStudy (anonymous):

please solve x^2-2x-10=0

OpenStudy (saifoo.khan):

u want the solutions?

OpenStudy (anonymous):

PLease

OpenStudy (saifoo.khan):

\[1-\sqrt{11}, 1+\sqrt{11}\]

OpenStudy (saifoo.khan):

have to submit it before midnight i guess?

OpenStudy (anonymous):

wow, that was quick! thx it is not -11?

OpenStudy (saifoo.khan):

what do u mean it's not -11?

OpenStudy (mimi_x3):

saif?

OpenStudy (saifoo.khan):

Mini? i will reply your email as soon as i wake up!

OpenStudy (mimi_x3):

You died?

OpenStudy (anonymous):

let a = 1, b = -2, c = -10; \[1\pm \sqrt{11}\]

OpenStudy (anonymous):

what i meant is that i got -11. was wrong. thx!

OpenStudy (saifoo.khan):

no, i went to bed to sleep, but was unable to sleep for like 2 hrs so came back. will go back any sec

OpenStudy (saifoo.khan):

ohh!

OpenStudy (mimi_x3):

oh goodnight